# Temperature dependent energy gap for Yu-Shiba-Rusinov states at the quantum phase transition

### v1

@@ Abstract
Motivated by recent experiments, which allow for fine tuning of the effective magnetic interaction between the impurity and the superconductor, we investigate the regime around the quantum phase transition where the system’s ground state changes from a weakly coupled free spin to a screened spin regime. At this transition we find that the YSR states remain at finite energies at low temperatures, thereby generating a gap in the spectrum, which is inconsistent with predictions of the original YSR theory. We investigate various gap-generating scenarios and determine that the local suppression of the order parameter, only captured by self-consistent calculations, generates the gap.
@@ I Introduction
Superconductivity and magnetism are generally considered antagonistic competing orders in condensed matter. In particular, if global magnetic interactions are introduced into a conventional superconductor, superconductivity will be suppressed and eventually vanishes as the magnetic coupling is increased. However, if the magnetic interaction is introduced locally by a single, isolated magnetic impurity, a new type of sub-gap impurity stat is introduced into the system. These Bogoliubov quasiparticle states were first predicted by Yu, Shiba and Rusinov [1, 2, 3, 4] and are commonly known as YSR states. Due to the particle-hole symmetry of the Bogoliubov spectrum, the YSR states always appear symmetrically around zero energy.
@@
In this work we capitalize on the recent experimental advancement on tuning YSR states and investigate in detail the behavior of the YSR states, at and around the QPT. We are particularly focusing on the possibility and physical reason for a finite energy gap at the QPT, such that the YSR states never reach zero energy. Such a gap in the energy spectrum at the QPT has been indicated in several earlier numerical works, but not much further elaborated upon. To the best of our knowledge, the first instances where a gap has been noted in the energy spectrum is Ref. 22, 23, 24, which modeled a magnetic impurity in the classical spin limit and relaxed the local order parameter of the superconducting substrate. A putative finite gap can also be found when closely investigating the results of later works using similar methods [25, 26]. Experimentally, a finite energy gap can also be envisaged in the data of Ref. [20], but, again, not further elaborated upon. Here we extend on these earlier indications of a finite gap and show both how and why a gap can appear in the YSR spectrum, using similar modeling to the afore-mentioned numerical works.
@@ II Modeling YSR states
Figure 1: a) Schematic picture of a superconducting substrate with indicated square lattice and with an absorbed magnetic impurity (red dot), treated within (left) the effective Anderson impurity model, with impurity on-site energy $\varepsilon_{i}$ and spin exchange coupling $J$ for the impurity, and with the impurity coupling to the substrate through the hopping amplitude $t_{0}$ and (right) classical Kondo model with effective spin exchange coupling $\tilde{J}$ and additional onsite energy of $\tilde{\varepsilon}_{i}$, both induced directly in the substrate. b) Ideal YSR state energies $\epsilon_{YSR}$ with states labeled according to their spin-polarization relative to the impurity spin, where superscript indicates occupied (--) and unoccupied ($+$) at zero temperature. The critical coupling $J_{C}$ marks the QPT.
@@
In general, both impurity models described above, quantitatively describe the same physics, also highlighted by the fact that they are linked via a Schrieffer-Wolff transformation [34] in the strong coupling limit. This transformation combines the effects of the coupling between the impurity and the substrate with the energy separation of the impurity levels, into the single scattering amplitude $\tilde{J}$. In this work we consider the mean-field Anderson impurity model when discussing general properties of YSR states and possibilities for an energy gap in the following Section III, and then we revert to the slightly simpler Kondo limit for the numerically demanding self-consistent calculations in Section IV. Overall, we find our results and conclusions to be fully consistent between the two models.
Before proceeding we note that the impurity treatment in this work, as in many others, do not incorporate possible quantum impurity behavior, e.g. see review [30] and references therein or for a comparison between different models [35, 36, 37]. Such a quantum impurity requires a many-body treatment, which requires computationally very expensive numerical methods, such as numerical renormalization group [38] or quantum Monte Carlo methods [39, 40, 41].
With our main results being based on a careful self-consistent treatment of the local order parameter, which also requires high resolution with respect to the magnetic coupling around the QPT, we are computationally limited to consider the classical spin limit. However, we note that an extension to quantum impurities is in principle feasible within an NRG treatment [42]. Importantly, the single-particle impurity models we use in our work have been shown to reproduce important features of the full quantum impurity problem [43, 36, 44, 45, 37, 46]. Moreover, for magnetic impurities with high spin, a classical Kondo treatment is also analytically motivated.
@@ II.1 YSR energies
Here we use the superscript 00 to indicate that this results holds in the idealized YSR limit of $J\gg\Delta$ and for the terms used in Eqn. (1-3).
@@ II.2 YSR gap possibilities
The second scenario to produce a gap in the spectrum is to in some indirect way couple the two diagonal blocks of the local Green’s function in Eq. (6). Even if there is no direct coupling present in this second scenario, it may still be possible to provide a change of the ingoing parameters that affects both blocks, such that the crossing at $J_{C}$ is no longer smooth, or even continuous, and, as a result, a gap may be generated. We illustrate schematically such a behavior in Fig. 2b), where there is a jump in the energy levels for both the upper and lower YSR states, such that no state actually cross zero energy. Obviously this scenario requires a dramatic modification to the Green’s function exactly at $J_{C}$ and may not be what is at first expected by just continuously tuning a single parameter $J$ for a single impurity in a thermodynamically large system. However, it is already well-established that the superconducting order parameter $\Delta(\textbf{x}_{0})$ at the impurity site $\textbf{x}_{0}$, is heavily suppressed and even shift non-continuously to a negative value at the QPT, denoted as a $\pi$-shift in the order parameter [22, 23, 30, 25, 26, 31, 32, 33]. Since the order parameter in turn determines the YSR energy levels through $\mathcal{H}$, such drastic change of the order parameter may be a candidate to also induce a discontinuous behavior of the energy levels at $J_{C}$, akin to Fig. 2b), resulting in a gap in the spectrum. A hint of such gap is in fact present already in early numerical data [22] as well as later works [24, 23, 25, 26], but has to our knowledge not been further discussed in the literature. We finally note that in this second scenario there is no mixing between the two YSR states and the spin polarization of each state therefore stays constant. We show in Section IV that this scenario of opening a gap can occur along with a strong temperature dependence.
@@ III Gap from interactions between YSR states
[…]
@@ IV Gap from local suppression of superconductivity
IV.1 Self-consistent modelling
To be able to perform accurate numerical modeling, we, for this part, resort to the classical Kondo limit of the magnetic impurity, which reduces the numerical complexity while still capturing the same essence of the magnetic impurity. To avoid any misunderstandings, we will in this Section use variables with a tilde, such as $\tilde{J}$, to highlight the use of the Kondo model, instead of the Anderson impurity model in the previous Section. As such, we consider a conventional spin-singlet $s$-wave superconductor substrate where a single site in the substrate hosts a magnetic impurity modeled by adding to that site the interaction $\tilde{J}\textbf{S}\textbf{s}_{\textbf{x}_{0}}$, where S is the impurity spin, $\textbf{s}_{\textbf{x}_{0}}$ is the local spin of the substrate’s electronic system at the impurity site $\textbf{x}_{0}$, and $\tilde{J}$ the coupling strength between the two. We then treat the magnetic impurity in the classical spin approximation, where $\textbf{S}\rightarrow\infty$ and $\tilde{J}\rightarrow 0$, such that the product $\tilde{J}\textbf{S}\rightarrow\tilde{\textbf{J}}$ remains finite. This approximation is especially suitable for large impurity spins or for systems with a strong anisotropy that aligns the spin in a particular direction [37] and, importantly for our purposes, the model is known to fundamentally capture the influence of a magnetic impurity on the superconductor [22, 25, 26, 32]. As a consequence, the impurity Hamiltonian is now given by
~$\tilde{\mathcal{H}}_{i}=\sum_{\sigma,\sigma^{\prime}}\left(\tilde{\textbf{J}}\left(\vec{\sigma}\right)_{\sigma\sigma^{\prime}}+\tilde{\varepsilon}_{i}\left(\sigma_{0}\right)_{\sigma\sigma^{\prime}}\right)c^{\dagger}_{\textbf{x}_{0}\sigma}c_{\textbf{x}_{0}\sigma^{\prime}},$ (9)
@@ IV.1 Self-consistent modelling
[…]
@@
where $\epsilon_{\nu}$ denotes the $\nu$-th eigenvalue of the full Hamiltonian ~=~$\tilde{\mathcal{H}}=\tilde{\mathcal{H}}_{i}+\tilde{\mathcal{H}}_{sc}$ and $u_{i\sigma}$, $v_{i\sigma}$ are the corresponding eigenstates at lattice site $i$ for spin $\sigma=\uparrow,\downarrow$. Further, $V_{sc}$ is the effective attractive interaction giving rise to superconductivity, effectively encoding for the attractive (assumed electron-phonon) interaction. We make the reasonable assumption that $V_{sc}$ is constant in the whole sample and not influenced by the single impurity. The Dirac-Fermi function $f_{T}(\epsilon)$ generates a temperature $T$ dependence, which we express for convenience in units of $k_{B}$.
We generate a self-consistent local order parameter $\Delta_{\textbf{i}}$ iteratively, by first diagonalizing $\tilde{\mathcal{H}}$, then calculating (12), and finally we use the new $\Delta_{\textbf{i}}$ in $\tilde{\mathcal{H}}$ to for next iteration. As convergence criteria, we accept a solution for $\Delta_{\textbf{i}}$ if the smallest difference between two iteration steps for any site is below $5\cdot 10^{-6}\,\Delta_{0}$, where $\Delta_{0}$ is the order parameter of the unperturbed substrate. Unless otherwise specified, we employ an attractive potential of $V_{sc}=1.825\,t$, leading to $\Delta_{0}\approx 0.24\,t$ and a superconducting transition temperature of $T_{C}\approx 0.14\,t$. For key results we also use other $V_{sc}$ to check the dependence on $V_{sc}$. We further limit our results to temperatures well below $T_{C}$ where the temperature effect on $\Delta_{0}$ itself is negligible. We use system sizes up to $1111\times 1$ 111 and, find, for the parameters mentioned above, the system to be adequately converged for sizes already of size $61\times 61$, which is the size we use for our numerical calculations unless otherwise indicated. To avoid effects from spurious edge states we apply periodic boundary conditions.
@@ IV.2 Energy spectrum
We start by discussing the zero temperature results in Fig. 3a). We find that the YSR states are approximately linearly decreasing (increasing) as a function of $\tilde{J}$ for $\tilde{J}$ below (above) $\tilde{J}_{C}$. This is fully consistent with the non-self-consistent results around the QPT in Fig. 1b). However, when increasing $\tilde{J}$ and when the states reach zero energy, they discontinuously jump to a finite energy just past the QPT. Thus, there is a distinct energy jump at $\tilde{J}_{C}$, albeit, we note technically no gap at zero energy at $T=0$. We can attribute this energy jump to an equally sudden suppression of the local order parameter $\Delta_{\textbf{x}_{0}}$ at $\tilde{J}_{C}$, including also a local $\pi$-shift, as shown in Fig. 3c) where we plot the local order parameter as a function of $\tilde{J}$ for the same temperatures. The discontinuous drop in the local order parameter at the impurity site can in turn be directly attributed to the $\ket{\downarrow}^{+}$ state suddenly becoming occupied for $\tilde{J}>\tilde{J}_{C}$ instead of the previously occupied $\ket{\uparrow}^{-}$ state. It has previously been established that the contribution of the $\ket{\downarrow}^{+}$ to $\Delta_{\textbf{x}_{0}}$ in the self-consistent gap equation Eq. (12) is out-of-resonance with the rest of the condensate, or equivalently, $\pi$-shifted compared to other states occupied states [32]. Hence, suddenly transferring occupation from the $\ket{\uparrow}^{-}$ state to the $\ket{\downarrow}^{+}$ state at $\tilde{J}_{C}$, the local order parameter is equally suddenly suppressed. This sudden suppression is the reason behind the jump in the YSR energies at $\tilde{J}_{C}$. We note that $\pi$-shift in the local order parameter is further energetically favored at the QPT, as it lowers the free energy of the system $F(\tilde{J})$. We discuss this further in the Appendix B. While not evident in Fig. 3a), we further observed that $\tilde{J}_{C}$ is shifted to a lower value compared to the constant $\Delta$ approximation, which is in line with previously reported results [22, 24, 23]. To summarize, the discontinuous jump of the superconducting order parameter and the discontinuous of the YSR energies, both at $\tilde{J}_{C}$, is intimately connected through the self-consistency gap equation Eq. (12).
At small, but finite temperatures we find the same linearly decreasing (increasing) YSR state energies for $\tilde{J}$ far below (above) $\tilde{J}_{C}$. In fact, the YSR state energies far below or above $\tilde{J}_{C}$ are insensitive to temperature. However, close to $\tilde{J}_{C}$ we find notable temperature effects. In the temperature regime $0<T<T^{*}$, with $T^{*}\approx 0.044\,T_{C}$ for the parameters used here, the YSR state energies do not cross through zero energy anymore but always remain at a finite value at $\tilde{J}_{C}$. The YSR energy levels still present similarities with the jump present at $T=0$, but the jump is decreasing with increasing temperature such that it is entirely smoothed out for temperatures above $T^{*}$. When the jump decreases, the gap instead increases for increasing temperatures. In Fig. 3b) we additionally plot a zoom-in around the QPT to display the finite temperature behavior with each data point corresponding to a separate calculation with different $\tilde{J}$. The figure clearly shows both the gap and the jump at $\tilde{J}_{C}$ is not an effect of too sparse sampling along the $\tilde{J}$-axis. Additionally, in color we show the magnetization of the two YSR states $\left\langle s_{Z}\right\rangle_{YSR}$ to highlight that also the character of the occupied YSR state changes abruptly at $\tilde{J}_{C}$. We can understand the behavior of both the jump and the gap by recognizing that finite temperature causes a finite occupation of the $\ket{\downarrow}^{+}$ state due to thermal fluctuations even at $\tilde{J}<\tilde{J}_{C}(T=0)$ (and likewise, not full occupation of the $\ket{\uparrow}^{-}$ state). As a consequence, the suppression of the local order parameter and its accompanied $\pi$-shift occurs already at lower $\tilde{J}_{C}$ compared to at $T=0$. This in turn leads to the QPT occurring at a lower $\tilde{J}_{C}$, i.e. $\tilde{J}_{C}(T>0)<\tilde{J}_{C}(T=0)$, as clearly seen in Fig. 3a). This move of $\tilde{J}_{C}$ further needs to be reconciled with the fact that the effects of thermal occupation is only a very low-energy phenomena, and therefore the YSR energies are still at the same energies for all temperatures away from the immediate neighborhood of the QPT. To accommodate both of these behaviors, with minimal other changes to the spectrum, the energy jump at $\tilde{J}_{C}$ becomes smaller at finite temperatures, while at the same time a gap necessarily develops at the QPT. For the parameters chosen in Fig. 3, the energy jump is always to more negative energies for the occupied state when increasing $J$ past $\tilde{J}_{C}$ (and inversely for the unoccupied state). However, this jump in energy at $\tilde{J}_{C}$ depends on the trend of the energy spectrum before and after the QPT and therefore, for other parameters this jump in energy can change direction, but is still aways present. To summarize, the low-temperature behavior is driven by thermal occupation, which decreases $\tilde{J}_{C}$ with increasing temperature, which in turn forces the spectrum to reduce the energy jump and instead develop an energy gap at $\tilde{J}_{C}$.
@@
Finally, for temperatures much above $T^{*}$, the YSR state energies again change smoothly through zero energy at $\tilde{J}_{C}$, similar to the constant $\Delta$ approximation and hence no effect of self-consistency is observed for these higher temperatures. Also, $\tilde{J}_{C}$ does not change with temperature in this high-temperature regime. Furthermore, the local order parameter $\Delta_{\textbf{x}_{0}}$ does not exhibit a jump at $\tilde{J}_{C}$, but instead has a constant positive value. As the local order parameter $\Delta_{\textbf{x}_{0}}$ is positive at $\tilde{J}_{C}$, this means, in turn, that the $\pi$-shift is shifted towards magnetic couplings $\tilde{J}$ larger than $\tilde{J}_{C}$, as also seen in Fig. 3c). Thus, the existence of the YSR energy gap is not directly related to the $\pi$-shift of the order parameter. The gap is instead tied to the sudden jump in the order parameter, which when it disappears, the YSR gap also disappears, while the $\pi$-shift still remain, albeit its onset is beyond $\tilde{J}_{C}$.
To summarize, the YSR state energies show a complex temperature behavior at $\tilde{J}_{C}$: a sharp energy jump at $T=0$ develops partially into an energy gap at finite temperatures, until at some temperature scale both jump and gap disappear, leaving a linear zero-energy crossing. All this behavior is intimately tied to the jump in the local superconducting order parameter at $\tilde{J}_{C}$. In Fig. 3d) we extract the minimum YSR state energies, i.e. the YSR state energy gap, as a function of temperature, which shows a notable energy gap for a range of small but finite temperatures.

### v2

@@ Abstract
Motivated by recent experiments, which allow for fine tuning of the effective magnetic interaction between the impurity and the superconductor, we investigate the regime around the quantum phase transition where the system’s ground state changes from a weakly coupled free spin to a screened spin regime. At this transition we find that the Yu-Shiba-Rusinov (YSR) states remain at finite energies at low temperatures, thereby generating a gap in the spectrum, which is inconsistent with predictions of the original YSR theory. We investigate various gap-generating scenarios and determine that the local suppression of the order parameter, only captured by self-consistent calculations, generates the gap.
@@ I Introduction
Superconductivity and magnetism are generally considered antagonistic competing orders in condensed matter. In particular, if global magnetic interactions are introduced into a conventional superconductor, superconductivity will be suppressed and eventually vanishes as the magnetic coupling is increased. However, if the magnetic interaction is introduced locally by a single, isolated magnetic impurity, a new type of sub-gap impurity state is introduced into the system. These Bogoliubov quasiparticle states were first predicted by Yu, Shiba and Rusinov [1, 2, 3, 4] and are commonly known as YSR states. Due to the particle-hole symmetry of the Bogoliubov spectrum, the YSR states always appear symmetrically around zero energy.
@@
In this work we capitalize on the recent experimental advancement on tuning YSR states and investigate in detail the behavior of the YSR states, at and around the QPT. We are particularly focusing on the possibility and physical reason for a finite energy gap at the QPT, such that the YSR states never reach zero energy. Such a gap in the energy spectrum at the QPT has been indicated in several earlier numerical works, but not much further elaborated upon. To the best of our knowledge, the first instances where a gap has been noted in the energy spectrum is Ref. 22, 23, 24, which modeled a magnetic impurity in the classical spin limit and relaxed the local order parameter of the superconducting substrate. A putative finite gap can also be found when closely investigating the results of later works using similar methods [25, 26]. In another model comprising two coupled impurity spins, the presence of a gap has also been observed for specific parameter choices, but this effect does not account for an energy gap for a single impurity spin [27]. Experimentally, a finite energy gap has also been reported in Ref. [20] for a single magnetic impurity, but, again, the underlying mechanism was not substantially elaborated upon, as the gap was not the focus of that work. Here we extend on these earlier results on a finite energy gap for the YSR states and show both how and why a gap can appear in the YSR spectrum, using similar modeling to the afore-mentioned numerical works.
@@ II Modeling YSR states
Figure 1: a) Schematic picture of a superconducting substrate with indicated square lattice and with an absorbed magnetic impurity (red dot), treated within (left) the effective Anderson impurity model, with impurity on-site energy $\varepsilon_{i}$ and spin exchange coupling $J$ for the impurity, and with the impurity coupling to the substrate through the hopping amplitude $t_{0}$ and (right) classical Kondo model with effective spin exchange coupling $\tilde{J}$ and additional onsite energy of $\tilde{\varepsilon}_{i}$, both induced directly in the substrate. b) Ideal YSR state energies $\epsilon_{YSR}$ with states labeled according to their spin-polarization relative to the impurity spin, where superscript indicates occupied (-) and unoccupied (++) at zero temperature. The critical coupling $J_{C}$ marks the QPT.
@@
In general, both impurity models described above, quantitatively describe the same physics, also highlighted by the fact that they are directly linked via a Schrieffer-Wolff transformation [35] in the strong coupling limit. This transformation combines the effects of the coupling between the impurity and the substrate with the energy separation of the impurity levels, into the single scattering amplitude $\tilde{J}$. In this work we consider the mean-field Anderson impurity model when discussing general properties of YSR states and possibilities for an energy gap in the following Section III, but then we need to revert to the slightly simpler Kondo limit for the numerically demanding self-consistent calculations in Section IV. Overall, we find our results and conclusions to be fully consistent between the two models, as also guaranteed by their direct link [35]. This involves numerical verification of key results between the two models.
Before proceeding we note that the impurity treatment in this work, as in many others, do not incorporate possible quantum impurity behavior, e.g. see review [31] and references therein or for a comparison between different models [36, 37, 38]. Such a quantum impurity requires a many-body treatment, which demands computationally very expensive numerical methods, such as numerical renormalization group (NRG) [39] or quantum Monte Carlo methods [40, 41, 42]. Still, the single-particle impurity models, or classical spin models, we use in our work been shown to reproduce important features of the full quantum impurity problem [36, 37, 43, 44, 38, 45]. Most importantly, classical spin models are well motivated both for magnetic impurities with high spin or for systems with axial magnetic anisotropy [38], and has also been shown to produce results in good agreement with recent experimental findings also in spin-1/2 systems [15, 20]. Finally, since our main results are based on a careful self-consistent treatment of the local order parameter, which also requires high resolution with respect to the magnetic coupling around the QPT, we are computationally limited to consider the classical spin limit. However, we note that an extension to quantum impurities is in principle feasible within an NRG treatment [46], albeit very computationally expensive.
@@ II.1 YSR energies
Here we use the superscript 0 to indicate that this results holds in the idealized YSR limit of $J\gg\Delta$ and for the terms used in Eqn. (1-3).
@@ II.2 YSR gap possibilities
The second scenario to produce a gap in the spectrum is to in some indirect way couple the two diagonal blocks of the local Green’s function in Eq. (6). Even if there is no direct coupling present in this second scenario, it may still be possible to provide a change of the ingoing parameters that affects both blocks, such that the crossing at $J_{C}$ is no longer smooth, or even continuous, and, as a result, a gap may be generated. We illustrate schematically such a behavior in Fig. 2b), where there is a jump in the energy levels for both the upper and lower YSR states, such that no state actually cross zero energy. Obviously this scenario requires a dramatic modification to the Green’s function exactly at $J_{C}$ and may not be what is at first expected by just continuously tuning a single parameter $J$ for a single impurity in a thermodynamically large system. However, it is already well-established that the superconducting order parameter $\Delta(\textbf{x}_{0})$ at the impurity site $\textbf{x}_{0}$, is heavily suppressed and even shift non-continuously to a negative value at the QPT, denoted as a $\pi$-shift in the order parameter [22, 23, 48, 31, 25, 26, 32, 33, 34]. Since the order parameter in turn determines the YSR energy levels through $\mathcal{H}$, such drastic change of the order parameter may be a candidate to also induce a discontinuous behavior of the energy levels at $J_{C}$, akin to Fig. 2b), resulting in a gap in the spectrum. A hint of such gap is in fact present already in early numerical data [22] as well as later works [24, 23, 48, 25, 26], but has to our knowledge not been further discussed in the literature. We finally note that in this second scenario there is no mixing between the two YSR states and the spin polarization of each state therefore stays constant. We show in Section IV that this scenario of opening a gap can occur along with a strong temperature dependence.
@@ III Gap from interactions between YSR states
Based on the above results, we conclude that for ideal YSR states, none of the additional terms in Eq. (8) produce a gap. However, there may still exist altered situations where the summation over k does not lead to a full cancellation as it requires inversion symmetry in the substrate plane. This can for example be the case if the superconductor is found to be very dirty, where we cannot necessarily expect inversion symmetry to prevail, not even on an average level, or, if the coupling between the impurity and the substrate is not fully localized (i.e. it has a k dependence) and additionally violates the inversion symmetry between k and $\textbf{k}$. In these cases the summation over k in Eq. (3) may not cause a full cancellation for some of the additional coupling terms in Eq. (8). For the sake of completeness, we report the results for creating an energy gap for such altered situations in Appendix A. We further point out that the induced magnitude of the energy gap at the QPT in these situations is generally temperature independent, unless the coefficients $\alpha_{i\>\!j}$ happens to have an explicit temperature dependency. This in sharp contrast to the gap that is caused from the indirect coupling through the self-consistently calculated local order parameter, that we discuss in the next Section (IV).
@@ IV Gap from local suppression of superconductivity
IV.1 Self-consistent modeling
To be able to perform accurate numerical modeling, we, for this part, resort to the classical Kondo limit of the magnetic impurity, which reduces the numerical complexity while still capturing the same essence of the magnetic impurity. To avoid any misunderstandings, we will in this Section use variables with a tilde, such as $\tilde{J}$, to highlight the use of the Kondo model, instead of the Anderson impurity model in the previous Section. As such, we consider a conventional spin-singlet $s$-wave superconductor substrate where a single site in the substrate hosts a magnetic impurity modeled by adding to that site the interaction $\tilde{J}\textbf{S}\textbf{s}_{\textbf{x}_{0}}$, where S is the impurity spin, $\textbf{s}_{\textbf{x}_{0}}$ is the local spin of the substrate’s electronic system at the impurity site $\textbf{x}_{0}$, and $\tilde{J}$ the coupling strength between the two. We then treat the magnetic impurity in the classical spin approximation, where $\textbf{S}\rightarrow\infty$ and $\tilde{J}\rightarrow 0$, such that the product $\tilde{J}\textbf{S}\rightarrow\tilde{\textbf{J}}$ remains finite. This approximation is especially suitable for large impurity spins or for systems with a strong anisotropy that aligns the spin in a particular direction [38] and, importantly for our purposes, the model is known to fundamentally capture the influence of a magnetic impurity on the superconductor [22, 48, 25, 26, 33]. As a consequence, the impurity Hamiltonian is now given by
$\tilde{\mathcal{H}}_{i}=\sum_{\sigma,\sigma^{\prime}}\left(\tilde{\textbf{J}}\left(\vec{\sigma}\right)_{\sigma\sigma^{\prime}}+\tilde{\varepsilon}_{i}\left(\sigma_{0}\right)_{\sigma\sigma^{\prime}}\right)c^{\dagger}_{\textbf{x}_{0}\sigma}c_{\textbf{x}_{0}\sigma^{\prime}},$ (9)
@@ IV.1 Self-consistent modeling
$\tilde{\mathcal{H}}_{sc}=\mu\sum_{\textbf{i},\sigma}c^{\dagger}_{\textbf{i}\sigma}c_{\textbf{i}\sigma}-t\sum_{\langle\textbf{i},\textbf{j}\rangle,\sigma}c^{\dagger}_{\textbf{i}\sigma}c_{\textbf{j}\sigma}+\sum_{\textbf{i}}\Delta_{\textbf{i}}c_{\textbf{i}\uparrow}c_{\textbf{i}\downarrow}+\text{h.c.},$ (11)
where $t$ is the hopping integral between two neighboring sites $\langle\textbf{i},\textbf{j}\rangle$, which we also use as the unit of energy. Further, $\mu$ denotes the onsite energy, which we set to $\mu=0.5\,t$ to move the Fermi level away from the van Hove singularity at $\mu=0$, in order to model a generic band structure. We also perform checks for other values of $\mu$, including at the van Hove singularity at $\mu=0$ and for $\mu\approx 3.62\,t$ where the Fermi level becomes circular. We find quantitatively similar results, except for a different scaling of the energies related to the order parameter $\Delta_{\textbf{i}}$, which can be attributed to a change of the density of states (DOS) at the Fermi level. This is in line with previous work studying the effects of the van Hove singularity on YSR states, where the YSR energies were found to depend on the change in the zero-energy DOS [52, 53].
@@
where $\epsilon_{\nu}$ denotes the $\nu$-th eigenvalue of the full Hamiltonian $\tilde{\mathcal{H}}=\tilde{\mathcal{H}}_{i}+\tilde{\mathcal{H}}_{sc}$ and $u_{i\sigma}$, $v_{i\sigma}$ are the corresponding eigenstates at lattice site $i$ for spin $\sigma=\uparrow,\downarrow$. Further, $V_{sc}$ is the effective attractive interaction giving rise to superconductivity, effectively encoding for the attractive (assumed electron-phonon) interaction. We make the reasonable assumption that $V_{sc}$ is constant in the whole sample and not influenced by the single impurity. The Dirac-Fermi function $f_{T}(\epsilon)$ generates a temperature $T$ dependence, which we express for convenience in units of $k_{B}$.
We generate a self-consistent local order parameter $\Delta_{\textbf{i}}$ iteratively, by first diagonalizing $\tilde{\mathcal{H}}$, then calculating (12), and finally we use the new $\Delta_{\textbf{i}}$ in $\tilde{\mathcal{H}}$ to for next iteration. As convergence criteria, we accept a solution for $\Delta_{\textbf{i}}$ if the smallest difference between two iteration steps for any site is below $5\cdot 10^{-6}\,\Delta_{0}$, where $\Delta_{0}$ is the order parameter of the unperturbed substrate. Unless otherwise specified, we employ an attractive potential of $V_{sc}=1.825\,t$, leading to $\Delta_{0}\approx 0.24\,t$ and a superconducting transition temperature of $T_{C}\approx 0.14\,t$. For key results we also use other $V_{sc}$ to check the dependence on $V_{sc}$. We further limit our results to temperatures well below $T_{C}$ where the temperature effect on $\Delta_{0}$ itself is negligible. We use system sizes up to $111\times 111$ and, find, for the parameters mentioned above, the system to be adequately converged for sizes already of size $61\times 61$, which is the size we use for our numerical calculations unless otherwise indicated. To avoid effects from spurious edge states we apply periodic boundary conditions.
@@ IV.2 Energy spectrum
[…]
@@
Finally, for temperatures much above $T^{*}$, the YSR state energies again change smoothly through zero energy at $\tilde{J}_{C}$, similar to the constant $\Delta$ approximation and hence no effect of self-consistency is observed for these higher temperatures. Also, $\tilde{J}_{C}$ does not change with temperature in this high-temperature regime. Furthermore, the local order parameter $\Delta_{\textbf{x}_{0}}$ does not exhibit a jump at $\tilde{J}_{C}$, but instead has a constant positive value. As the local order parameter $\Delta_{\textbf{x}_{0}}$ is positive at $\tilde{J}_{C}$, this means, in turn, that the $\pi$-shift is moved towards magnetic couplings $\tilde{J}$ larger than $\tilde{J}_{C}$, as also seen in Fig. 3c). Thus, the existence of the YSR energy gap is not directly related to the $\pi$-shift of the order parameter. The gap is instead tied to the sudden jump in the order parameter, which when it disappears, the YSR gap also disappears, while the $\pi$-shift still remain, albeit its onset is beyond $\tilde{J}_{C}$.
Additionally, we find that, regardless of temperature, the local suppression of the order parameter does not notably affect the superconducting energy gap in the spectrum, with any change at least a magnitude smaller than the energy gap between the YSR states. Similar effects have also been discussed or shown in [22, 24, 23, 25].
To summarize, the YSR state energies show a complex temperature behavior at $\tilde{J}_{C}$: a sharp energy jump at $T=0$ develops partially into an energy gap at finite temperatures, until at some temperature scale both jump and gap disappear, leaving a linear zero-energy crossing. All this behavior is intimately tied to the jump in the local superconducting order parameter at $\tilde{J}_{C}$. In Fig. 3d) we extract the minimum YSR state energies, i.e. the YSR state energy gap, as a function of temperature, which shows a notable energy gap for a range of small, but finite temperatures. To cross-reference our results we also calculate the suppression of the local order parameter in the Anderson impurity model for some key temperatures and find that the $\pi$-shift at the QPT also there causes a finite gap in the YSR spectrum at low temperatures.